# Exploring Typographic Visual Prompts Injection Threats in Cross-Modality Generation Models

### v3

@@ 1 Introduction
Recently, with the rapid advancement of Artificial General Intelligence (AGI), various Generation Models (GMs) have achieved remarkable success in diverse cross-modality tasks. Due to the ubiquity and rich information of vision modality in the real world, Cross-Vision GMs, apable of handling Vision-Language Perception (VLP) and Image-to-Image (I2I) generation tasks, receive extensive attention. Correspondingly, Large Vision-Language Models (LVLMs) are primarily used for VLP tasks, while I2I GMs are designed for I2I generation. The typical architecture of LVLMs (Liu et al., 2024, 2023; Chen et al., 2024; Lu et al., 2024; Team, 2025; Wang et al., 2024) comprises a vision encoder, which shares the same structure as Vision-Language Models exemplified by CLIP (Radford et al., 2021a), integrated with various Large Language Models (LLMs) (Touvron et al., 2023; Gao et al., 2023). Current I2I GMs can be broadly categorized into two types: (1) CLIP-guided diffusion models (Ramesh et al., 2022; Ye et al., 2023; Rombach et al., 2022; Podell et al., 2023), which use the CLIP vision encoder to jointly perceive visual and textual information; (2) Multimodal Large Language Models (MLLMs)-based I2I GMs (OpenAI, 2025; ByteDance, 2025), which treat image generation as a modality-specific output task within the corresponding MLLMs.
@@
In previous studies (Cheng et al., 2024, 2025; Wang et al., 2025; Chung et al., 2024; Levy and Liebmann, 2024), typographic word injection demonstrates significant security threats to various Cross-Vision GMs.  (Cheng et al., 2024; Wang et al., 2025; Chung et al., 2024; Levy and Liebmann, 2024) reveal that injecting a simple typographic word into the input images of LVLMs would significantly distract the final language output in various VLP tasks. Simultaneously,  (Cheng et al., 2025) demonstrates that printing typographic words into the input of CLIP-guided Diffusion Models (DMs) causes the generated images to incorporate relevant semantic information from the injected words. Comprehensively analyzing the impact of typographic words on the performance of LVLMs and I2I GMs helps uncover a potential, yet widely unrecognized, security threat under the vision modality. Additionally, a threat known as visual prompt injection (Kimura et al., 2024; Gong et al., 2023; Clusmann et al., 2025; Zhang et al., 2024) could disrupt the final output of LVLMs by injecting visual prompts into the input images that are unrelated to the textual prompts in the language modality. Actually, compared to traditional typographic words, visual prompts could be regarded as a more sophisticated form of typography. And this visual prompt is proven to induce significant security vulnerabilities in various current VLP tasks across different domains.  Kimura et al. (2024); Gong et al. (2023); Zhang et al. (2025) demonstrate that visual prompts can incur larger threats in jailbreak tasks.  Clusmann et al. (2025) and Zhang et al. (2024) highlight the security issues arising from typographic visual prompts in oncology examinations and GUI-agent operations. However, to date, compared to the comprehensive characteristic analysis of typographic word attacks in Cross-Vision modality tasks (Cheng et al., 2024, 2025), the threat induced by visual prompts still requires systematic exploration.
@@
In this paper, we systematically analyze the threats posed by Typographic Visual Prompt Injection (TVPI) across various Cross-Vision GMs. Based on the dataset construction approach in (Cheng et al., 2024, 2025), we propose the TVPI Dataset. The TVPI Dataset offers VLP and I2I subtype datasets to facilitate TVPI threat evaluation on LVLMs and I2I-GMs. The dataset incorporates 4 and 2 tasks for TVP and I2I subtypes separately, each defined by different instruction prompts. The visual prompts used in the dataset are further categorized into three thematic groups, each containing two target semantic concepts. Each subtype Dataset contains selected Clean images for attack, the Factor Modification (FM) with varied visual prompt factors, and the Different Target Word (DTW) to verify the TVPI threat across diverse application scenarios. In addition, we introduce a dedicated subtype to assess the vulnerability of TVPI attacks on various closed-source commercial Cross-Vision GMs. Figure 1 illustrates the overall process of executing TVPI, and demonstrates that TVPI effectively causes open-source and closed-source Cross-Vision GMs (LVLMs and I2I GMs) to deviate from the target semantics in the visual prompt across 4 VLP tasks and 2 I2I tasks. Through the above explorations, we further deepen the understanding of TVPI threats in different Cross-Vision GMs. Our contributions are as follows:
@@ 2 Related Works
Generation Models   Large Vision-Language Models (LVLMs) integrate vision-language modality information to generate final language outputs. This evolution is marked by the integration of pre-trained vision encoders and large language models (LLMs), enabling LVLMs to process and generate language based on visual inputs. Recent advancements include architectures that employ learnable queries to distill visual information and align it with LLM-generated text, as well as models like LLaVA (Liu et al., 2024, 2023), InternVL (Chen et al., 2024), Ovis (Lu et al., 2024), and Qwen (Team, 2025; Wang et al., 2024), which use projection layers to bridge visual features and textual embeddings. Additionally, the commercial closed-source LVLMs Claude-3.5-Sonnet (Anthropic) (Anthropic, 2025) and GPT-4o (OpenAI) (OpenAI, 2025) garner significant attention in contemporary society due to their advanced capabilities and widespread applications. Concretely, the application of LVLMs in VLP tasks extends to scenarios such as medical diagnosis (Xia et al., 2024; Hu et al., 2024), business operations (Huang et al., 2023; Pan et al., 2024), and education (Cherian et al., 2024). For Image-to-Image (I2I) Generation Models, previous architectures such as GANs (Goodfellow et al., 2020), VAEs (Kingma, 2013), and their variants (Heusel et al., 2017; Kong et al., 2020) demonstrate performance to a certain extent. However, diffusion-based models, particularly DDPM (Ho et al., 2020) and its variants (Nair et al., 2023; Li et al., 2024), have gained prominence due to their superior performance. Among these, CLIP-guided diffusion models, such as DALL-E 2 (UnCLIP) (Ramesh et al., 2022) and IP-Adapter (Ye et al., 2023), integrate the CLIP vision encoder (Radford et al., 2021b) to enhance visual semantic perception, enabling the generation of highly realistic, diverse, and semantically rich images. These models have become dominant in both research and commercial applications. Concurrently, the development of Multimodal Large Language Models (MLLMs) like GPT-4 (OpenAI) (OpenAI, 2025) and Dreamina (ByteDance) (ByteDance, 2025). While I2I tasks can also be expanded to fields such as artistic creation (Zhang et al., 2023; Wang et al., 2023), fundamental scientific exploration (Bauer and Metzler, 2012; Leven and Levy, 2019), and historical archaeology (Jaramillo and Sipiran, 2024; Cardarelli, 2025).
@@
Typographic Threats    Cheng et al. (2024, 2025) comprehensively evaluate threats of typographic words in LVLMs and I2I GMs.  Wang et al. (2025); Chung et al. (2024); Levy and Liebmann (2024) provide deeper explorations of the vulnerability of typographic words across various domains. For threats incurred by Typographic Visual Prompt Injection,  Kimura et al. (2024); Gong et al. (2023); Zhang et al. (2025) demonstrates that in the jailbreak attack task on LVLMs, visual prompts initiated from the vision modality present a greater threat compared to text prompts from the language modality. The threats caused by TVPI are also certified to exist in real-world application scenarios, including oncology examinations (Clusmann et al., 2025) and GUI-agent operation (Zhang et al., 2024).
@@ 3.1 Typographic Visual Prompts Injection Dataset
Clean and Factor Modification (FM) Setting   The base Clean images of the VLP and I2I subtypes are divided into 2000 and 500 examples, respectively. For the VLP subtype Dataset, we conduct experiments across four distinct subtasks that require identifying different object attributes: category, color, quantity, and size. Specifically, for the category subtask, we select 500 images from the ImageNet (Deng et al., 2009), along with a fixed text prompt "What object is depicted in the image?" for each image. In the color subtask, we employ 500 images from Visual7W (Zhu et al., 2016) with diverse queries inquiring about object color within each image. For the quantity subtask, we utilize 500 images from TallyQA (Acharya et al., 2019) paired with varied queries regarding object quantity in each image. In the size subtask, we choose 500 images from MSCOCO (Lin et al., 2014), along with a fixed text prompt "What item takes up the most size in the image?" for each image.
In I2I subtype Dataset, we design two distinct subtasks: photographic style transfer and full-body pose generation. Each subtask evaluates different aspects of image-to-image generation capabilities. For photographic style transfer, we employ the text prompt "analog film photo, faded film, desaturated, 35mm photo" to transform source images into ones with an analog aesthetic. For the full-body pose generation subtask, we use the text prompt "a youthful figure on the stage, full body view, dynamic pose" to generate human figures in specified poses. For each subtask, we select 500 images from CelebA-HQ (Karras, 2017; Liu et al., 2018)
@@ 4.1 Experimental Setting
Models   For the Vision-Language Perception (VLP) task, we conduct extensive experiments on current advanced open-source Large Vision Language Models series LLaVA-v1.6 (Liu et al., 2024, 2023), InternVL-v2.5 (Chen et al., 2024), Ovis-v2 (Lu et al., 2024), and Qwen-v2.5-VL (Team, 2025; Wang et al., 2024). For closed-source LVLMs, we evaluate two widely-used commercial models with APIs: Claude-3.5-Sonnet (Anthropic) (Anthropic, 2025) and GPT-4o (OpenAI) (OpenAI, 2025). For the Image-to-Image (I2I) task, we conduct experiments across DALL-E 2 or UnCLIP (Ramesh et al., 2022) and IP-Adapter (Ye et al., 2023). For IP-Adapter, we adopt three popular diffusion models, which are Stable Diffusion v1.5 (SD1.5) (Rombach et al., 2022), Stable Diffusion XL (SDXL) (Podell et al., 2023), and FLUX.1-dev (FLUX) (Esser et al., 2024). For closed-source I2I GMs, we evaluate two popular models, GPT-4 (OpenAI) (OpenAI, 2025) and Dreamina (ByteDance) (ByteDance, 2025).
Datasets   We adopt Typographic Visual Prompt Injection (TVPI) Dataset. The VLP and I2I subtype datasets are used to evaluate the TVPI threats of various Cross-Vision GMs (LVLMs and I2I GMs) under different factors and attack targets. The closed-source subtype dataset is specifically designed to execute on commercial APIs and official websites (Anthropic, 2025; OpenAI, 2025; ByteDance, 2025) of various GMs.
@@
In the I2I task, we employ CLIPScore (Radford et al., 2021a) to measure semantic alignment between generated images and their corresponding inserted attack targets. Higher CLIPScore values indicate stronger semantic similarity between the generated image and attack targets, suggesting more significant influence from the typographic visual prompts. Additionally, we utilize Fréchet Inception Distance (FID) (Heusel et al., 2017) to quantify distribution differences between images generated from visual-prompt-injected inputs and their corresponding clean originals. Larger FID scores signify greater deviation from source images, demonstrating stronger attack impact.
@@ 4.3.2 Impact on Closed-Source Models
For the VLP task, Table 3 shows that Claude-3.5-Sonnet (Anthropic) (Anthropic, 2025) and GPT-4o (OpenAI) (OpenAI, 2025) are severely affected by typographic visual prompts. In the I2I task, as illustrated in Figure 2, the generated images from both GPT-4 (OpenAI) (OpenAI, 2025) and Dreamina (ByteDance) (ByteDance, 2025) exhibit clear influence from typographic visual prompts.
@@ 4.4.1 Comparison with Typographic Word Injection
We also compare the typographic visual prompt injection with the typographic word injection mentioned in the work (Cheng et al., 2024). Specifically, we reduce the typographic visual prompt to only the attack target word, constituting the typographic word injection. For the VLP task, Figure 3 demonstrates that typographic word has little impact on models’ output, while typographic visual prompts cause a high ASR. In the I2I task, Figure 2 shows that typographic word injection has less influence on the generated images from closed-source models GPT-4 and Dreamina, when compared to the effectiveness of typographic visual prompts.

### v4

@@ 1 Introduction
Recently, with the rapid advancement of Artificial General Intelligence (AGI), various Generation Models (GMs) have achieved remarkable success in diverse cross-modality tasks. Due to the ubiquity and rich information of vision modality in the real world, Cross-Vision GMs, apable of handling Vision-Language Perception (VLP) and Image-to-Image (I2I) generation tasks, receive extensive attention. Correspondingly, Large Vision-Language Models (LVLMs) are primarily used for VLP tasks, while I2I GMs are designed for I2I generation. The typical architecture of LVLMs (liu2024llavanext; liu2023improvedllava; chen2024expanding; lu2024ovis; qwen2.5-VL; Qwen2VL) comprises a vision encoder, which shares the same structure as Vision-Language Models exemplified by CLIP (radford2021learning), integrated with various Large Language Models (LLMs) (touvron2023llama; gao2023llama). Current I2I GMs can be broadly categorized into two types: (1) CLIP-guided diffusion models (ramesh2022hierarchical; ye2023ip; rombach2022high; podell2023sdxl), which use the CLIP vision encoder to jointly perceive visual and textual information; (2) Multimodal Large Language Models (MLLMs)-based I2I GMs (OpenAI2025; ByteDance2025), which treat image generation as a modality-specific output task within the corresponding MLLMs.
@@
In previous studies (cheng2024unveiling; cheng2024uncovering; wang2025typographic; chung2024transfer; levy2024deepfake), typographic word injection demonstrates significant security threats to various Cross-Vision GMs.  (cheng2024unveiling; wang2025typographic; chung2024transfer; levy2024deepfake) reveal that injecting a simple typographic word into the input images of LVLMs would significantly distract the final language output in various VLP tasks. Simultaneously,  (cheng2024uncovering) demonstrates that printing typographic words into the input of CLIP-guided Diffusion Models (DMs) causes the generated images to incorporate relevant semantic information from the injected words. Comprehensively analyzing the impact of typographic words on the performance of LVLMs and I2I GMs helps uncover a potential, yet widely unrecognized, security threat under the vision modality. Additionally, a threat known as visual prompt injection (kimura2024empirical; gong2023figstep; clusmann2025prompt; zhang2024attacking) could disrupt the final output of LVLMs by injecting visual prompts into the input images that are unrelated to the textual prompts in the language modality. Actually, compared to traditional typographic words, visual prompts could be regarded as a more sophisticated form of typography. And this visual prompt is proven to induce significant security vulnerabilities in various current VLP tasks across different domains.  kimura2024empirical; gong2023figstep; zhang2025fc demonstrate that visual prompts can incur larger threats in jailbreak tasks.  clusmann2025prompt and zhang2024attacking highlight the security issues arising from typographic visual prompts in oncology examinations and GUI-agent operations. However, to date, compared to the comprehensive characteristic analysis of typographic word attacks in Cross-Vision modality tasks (cheng2024unveiling; cheng2024uncovering), the threat induced by visual prompts still requires systematic exploration.
@@
In this paper, we systematically analyze the threats posed by Typographic Visual Prompt Injection (TVPI) across various Cross-Vision GMs. Based on the dataset construction approach in (cheng2024unveiling; cheng2024uncovering), we propose the TVPI Dataset. The TVPI Dataset offers VLP and I2I subtype datasets to facilitate TVPI threat evaluation on LVLMs and I2I-GMs. The dataset incorporates 4 and 2 tasks for TVP and I2I subtypes separately, each defined by different instruction prompts. The visual prompts used in the dataset are further categorized into three thematic groups, each containing two target semantic concepts. Each subtype Dataset contains selected Clean images for attack, the Factor Modification (FM) with varied visual prompt factors, and the Different Target Word (DTW) to verify the TVPI threat across diverse application scenarios. In addition, we introduce a dedicated subtype to assess the vulnerability of TVPI attacks on various closed-source commercial Cross-Vision GMs. Figure 1 illustrates the overall process of executing TVPI, and demonstrates that TVPI effectively causes open-source and closed-source Cross-Vision GMs (LVLMs and I2I GMs) to deviate from the target semantics in the visual prompt across 4 VLP tasks and 2 I2I tasks. Through the above explorations, we further deepen the understanding of TVPI threats in different Cross-Vision GMs. Our contributions are as follows:
@@ 2 Related Works
Generation Models   Large Vision-Language Models (LVLMs) integrate vision-language modality information to generate final language outputs. This evolution is marked by the integration of pre-trained vision encoders and large language models (LLMs), enabling LVLMs to process and generate language based on visual inputs. Recent advancements include architectures that employ learnable queries to distill visual information and align it with LLM-generated text, as well as models like LLaVA (liu2024llavanext; liu2023improvedllava), InternVL (chen2024expanding), Ovis (lu2024ovis), and Qwen (qwen2.5-VL; Qwen2VL), which use projection layers to bridge visual features and textual embeddings. Additionally, the commercial closed-source LVLMs Claude-3.5-Sonnet (Anthropic) (Anthropic2025) and GPT-4o (OpenAI) (OpenAI2025) garner significant attention in contemporary society due to their advanced capabilities and widespread applications. Concretely, the application of LVLMs in VLP tasks extends to scenarios such as medical diagnosis (xia2024cares; hu2024omnimedvqa), business operations (huang2023lvlms; pan2024flowlearn), and education (cherian2024evaluating). For Image-to-Image (I2I) Generation Models, previous architectures such as GANs (goodfellow2020generative), VAEs (kingma2013auto), and their variants (heusel2017gans; kong2020hifi) demonstrate performance to a certain extent. However, diffusion-based models, particularly DDPM (ho2020denoising) and its variants (nair2023ddpm; li2024spd), have gained prominence due to their superior performance. Among these, CLIP-guided diffusion models, such as DALL-E 2 (UnCLIP) (ramesh2022hierarchical) and IP-Adapter (ye2023ip), integrate the CLIP vision encoder (pmlr-v139-radFord21a) to enhance visual semantic perception, enabling the generation of highly realistic, diverse, and semantically rich images. These models have become dominant in both research and commercial applications. Concurrently, the development of Multimodal Large Language Models (MLLMs) like GPT-4 (OpenAI) (OpenAI2025) and Dreamina (ByteDance) (ByteDance2025). While I2I tasks can also be expanded to fields such as artistic creation (zhang2023inversion; wang2023stylediffusion), fundamental scientific exploration (bauer2012generalized; leven2019quantifying), and historical archaeology (jaramillo2024cultural; cardarelli2025pypotteryink).
@@
Typographic Threats    cheng2024unveiling; cheng2024uncovering comprehensively evaluate threats of typographic words in LVLMs and I2I GMs.  wang2025typographic; chung2024transfer; levy2024deepfake provide deeper explorations of the vulnerability of typographic words across various domains. For threats incurred by Typographic Visual Prompt Injection,  kimura2024empirical; gong2023figstep; zhang2025fc demonstrates that in the jailbreak attack task on LVLMs, visual prompts initiated from the vision modality present a greater threat compared to text prompts from the language modality. The threats caused by TVPI are also certified to exist in real-world application scenarios, including oncology examinations (clusmann2025prompt) and GUI-agent operation (zhang2024attacking).
@@ 3.1 Typographic Visual Prompts Injection Dataset
Clean and Factor Modification (FM) Setting   The base Clean images of the VLP and I2I subtypes are divided into 2000 and 500 examples, respectively. For the VLP subtype Dataset, we conduct experiments across four distinct subtasks that require identifying different object attributes: category, color, quantity, and size. Specifically, for the category subtask, we select 500 images from the ImageNet (deng2009imagenet), along with a fixed text prompt "What object is depicted in the image?" for each image. In the color subtask, we employ 500 images from Visual7W (zhu2016visual7w) with diverse queries inquiring about object color within each image. For the quantity subtask, we utilize 500 images from TallyQA (acharya2019tallyqa) paired with varied queries regarding object quantity in each image. In the size subtask, we choose 500 images from MSCOCO (lin2014microsoft), along with a fixed text prompt "What item takes up the most size in the image?" for each image.
In I2I subtype Dataset, we design two distinct subtasks: photographic style transfer and full-body pose generation. Each subtask evaluates different aspects of image-to-image generation capabilities. For photographic style transfer, we employ the text prompt "analog film photo, faded film, desaturated, 35mm photo" to transform source images into ones with an analog aesthetic. For the full-body pose generation subtask, we use the text prompt "a youthful figure on the stage, full body view, dynamic pose" to generate human figures in specified poses. For each subtask, we select 500 images from CelebA-HQ (karras2017progressive; liu2018large)
@@ 4.1 Experimental Setting
Models   For the Vision-Language Perception (VLP) task, we conduct extensive experiments on current advanced open-source Large Vision Language Models series LLaVA-v1.6 (liu2024llavanext; liu2023improvedllava), InternVL-v2.5 (chen2024expanding), Ovis-v2 (lu2024ovis), and Qwen-v2.5-VL (qwen2.5-VL; Qwen2VL). For closed-source LVLMs, we evaluate two widely-used commercial models with APIs: Claude-3.5-Sonnet (Anthropic) (Anthropic2025) and GPT-4o (OpenAI) (OpenAI2025). For the Image-to-Image (I2I) task, we conduct experiments across DALL-E 2 or UnCLIP (ramesh2022hierarchical) and IP-Adapter (ye2023ip). For IP-Adapter, we adopt three popular diffusion models, which are Stable Diffusion v1.5 (SD1.5) (rombach2022high), Stable Diffusion XL (SDXL) (podell2023sdxl), and FLUX.1-dev (FLUX) (esser2024scaling). For closed-source I2I GMs, we evaluate two popular models, GPT-4 (OpenAI) (OpenAI2025) and Dreamina (ByteDance) (ByteDance2025).
Datasets   We adopt Typographic Visual Prompt Injection (TVPI) Dataset. The VLP and I2I subtype datasets are used to evaluate the TVPI threats of various Cross-Vision GMs (LVLMs and I2I GMs) under different factors and attack targets. The closed-source subtype dataset is specifically designed to execute on commercial APIs and official websites (Anthropic2025; OpenAI2025; ByteDance2025) of various GMs.
@@
In the I2I task, we employ CLIPScore (radford2021learning) to measure semantic alignment between generated images and their corresponding inserted attack targets. Higher CLIPScore values indicate stronger semantic similarity between the generated image and attack targets, suggesting more significant influence from the typographic visual prompts. Additionally, we utilize Fréchet Inception Distance (FID) (heusel2017gans) to quantify distribution differences between images generated from visual-prompt-injected inputs and their corresponding clean originals. Larger FID scores signify greater deviation from source images, demonstrating stronger attack impact.
@@ 4.3.2 Impact on Closed-Source Models
For the VLP task, Table 3 shows that Claude-3.5-Sonnet (Anthropic) (Anthropic2025) and GPT-4o (OpenAI) (OpenAI2025) are severely affected by typographic visual prompts. In the I2I task, as illustrated in Figure 2, the generated images from both GPT-4 (OpenAI) (OpenAI2025) and Dreamina (ByteDance) (ByteDance2025) exhibit clear influence from typographic visual prompts.
@@ 4.4.1 Comparison with Typographic Word Injection
We also compare the typographic visual prompt injection with the typographic word injection mentioned in the work (cheng2024unveiling). Specifically, we reduce the typographic visual prompt to only the attack target word, constituting the typographic word injection. For the VLP task, Figure 3 demonstrates that typographic word has little impact on models’ output, while typographic visual prompts cause a high ASR. In the I2I task, Figure 2 shows that typographic word injection has less influence on the generated images from closed-source models GPT-4 and Dreamina, when compared to the effectiveness of typographic visual prompts.